Alexa and Zack are solving the following problem: The number of miles on Car A is 50 miles more than the number of miles on Car B, and the number of miles on Car B is 30 miles more than the number of miles on Car C. All the cars travel 50 miles in 1 hour. After 1 hour, twice the number of miles Car A is 70 miles less than 3 times the number of miles on Car C. How many miles were there on Car B initially? Alexa assumes there are m miles on Car B. Zack assumes there are m miles on Car C. Will Zack's answer be the same as Alexa's answer? Explain.
Found 3 solutions by ankor@dixie-net.com, Silver_surfer, greenestamps:
Answer by ankor@dixie-net.com(22740)

Alexa and Zack are solving the following problem:
The number of miles on Car A is 50 miles more than the number of miles on Car B,
a = b+50
and the number of miles on Car B is 30 miles more than the number of miles on Car C.
b = c+30
or
c = b-30
All the cars travel 50 miles in 1 hour.
All travel at 50 mph
After 1 hour, twice the number of miles Car A is 70 miles less than 3 times the number of miles on Car C.
2(a+50) = 3(c+50) - 70
2a + 100 = 3c + 150 - 70
2a = 3c + 150 - 70 - 100
2a = 3c - 20
Replace a with (b+50), replace c with (b-30)
2(b+50) = 3(b-30) - 20
2b + 100 = 3b - 90 - 20
2b + 100 = 3b - 110
2b = 3b - 110 - 100
2b = 3b - 210
210 = 3b - 2b
b = 210 mi car B initially
then
a = 210 + 50
a = 260 car A initially
and
c = 210 - 30
c = 180 mi Car C initially

Alexa assumes there are m miles on Car B.\
Zack assumes there are m miles on Car C.
Will Zack's answer be the same as Alexa's No

Answer by Silver_surfer(1)

Alexa and Zack are solving the following problem: The number of miles on Car A is 50 miles more than the number of miles on Car B, and the number of miles on Car B is 30 miles more Car C. All the cars travel 50 miles in 1 hour. After 1 hour, twice the number of miles on Car A is 70 miles less than 3 times the number of miles on Car C.
How many miles were there on Car B initially?
================================================================================
set Car A = a
set Car B = b
set Car C = c

a = 50 + b
b = 30 + c
2a = 3c - 70 rewritten c = (2a+70)/3

b = 30 + c

b = 30 + (2a+70)/3 substitute c

b = 30 + [2(50 + b) + 70 )]/3 substitute a

3b = 90 + 100 + 2b + 70 multiply both sides by 3

b = 260 simplify

Car B had 260 miles initially.
Car A had 310 miles
Car C had 230 miles

================================================================================

Alex assumes Car B has m miles.
Zack assumes Car C has m miles.

Will Zack’s answer be the same as Alex’s? No.

There can be one correct answer and since Car B and Car C had different miles, Zack is incorrect.

Answer by greenestamps(13195)



Ignore the solution from the other tutor. It does not use all of the given information to set up the problem correctly; and it does not interpret the question that is asked correctly.

The two students are solving the same problem in different ways; as long as their work is correct, they should (and will) get the same answer.

The given information is that car A has 50 more miles than car B and car B has 30 more miles than car C.

Alexa's solution....

Let m = miles on car B
Then m+50 = miles on car A
and m-30 = miles on car C

After 1 hour at 50mph, the numbers of miles on each car are
A: m+100
B: m+50
C: m+20

At that time, twice the number of miles on car A is 70 less than 3 times the number of miles on car C:

2(m+100)=3(m+20)-70
2m+200=3m-10
m=210

Alexa's answer: the number of miles on car B initially was m=210.

Zack's solution....

Let m = miles on car C
Then m+30 = miles on car B
and m+80 = miles on car A

After 1 hour at 50mph, the numbers of miles on each car are
A: m+130
B: m+80
C: m+50

At that time, twice the number of miles on car A is 70 less than 3 times the number of miles on car C:

2(m+130)=3(m+50)-70
2m+260=3m+80
m=180

Zack's answer: the number of miles on car B initially was m+30=210.

ANSWER: Zack's answer will be the same as Alexa's.